The number 3 (five) (written in base five) is equivalent to what symbol in base two.
Answer by khwang(438)

3(five) = 3 in decimal
3 = 2*1 + 1 ( remainder 1)
1 = 2*0 + 1 ( remainder 1)
So, 3 = 11 (two)

Kenny